OpenStudy (anonymous):

simplify the expression (a^2)^-3(a^3b)^2(b^3)^4

OpenStudy (anonymous):

Solve for b over the real numbers: b^14 = 0 Eliminate the exponent. Take the 14^th root of both sides: Answer: | | b = 0

OpenStudy (anonymous):

how do i do that

OpenStudy (debbieg):

Is this the expression? \(\Large (a^2)^{-3}(a^3b)^2(b^3)^4\) There is nothing to "solve for b", it isn't an equation... it's just an expression to be simplified. But is what I have above, correct? It's hard to tell from plain text sometimes. :)

OpenStudy (anonymous):

yes

OpenStudy (debbieg):

You'll just need a few of the rules for exponents: \(\Large (x^m)^{n}=x^{mn}\) \(\Large (xy)^m=x^{m}y^m\) \(\Large x^{-1}=\dfrac{1}{x} \) \(\Large x^{-m}=\dfrac{1}{x^m} \)

OpenStudy (anonymous):

ok

OpenStudy (debbieg):

So start with the exponents outside each set of ( ) and simplify that.... \(\Large (a^2)^{-3}(a^3b)^2(b^3)^4\\ \Large =a^{\Box} (a^{\Box}b^{\Box})(b^{\Box})\) What goes in each "box"

OpenStudy (anonymous):

im confused

OpenStudy (debbieg):

Oh, I forgot one other rule you are going to need: \(\Large x^m\cdot x^n=x^{m+n}\)

OpenStudy (debbieg):

OK, tell me where you're confused. Those rules for exponents that I posted above, you've seen those before, right? Do you understand them?

OpenStudy (anonymous):

the box part that confused me

OpenStudy (debbieg):

\(\Large (x^m)^{n}=x^{mn}\) so, e.g., \(\Large (3^2)^{5}=3^{10}\) \(\Large (z^2)^{-3}=z^{-6}\) etc.

OpenStudy (anonymous):

oh ok i understand that

OpenStudy (debbieg):

ah, ok.. sorry, I was just trying to show you what you needed to do first. One step at a time... you have: \(\Large (a^2)^{-3}(a^3b)^2(b^3)^4\) So tell me what you can do with \(\Large (a^2)^{-3}\) ?? What is \(\Large (a^2)^{-3}=\)?

OpenStudy (anonymous):

\[a ^{-6}\]

OpenStudy (debbieg):

Good! Now how about \(\Large (a^3b)^2=?\)

OpenStudy (anonymous):

\[a^{6}b ^{2}\]

OpenStudy (debbieg):

Perfect. \(\Large (b^3)^4=?\)

OpenStudy (anonymous):

\[b ^{12}\]

OpenStudy (debbieg):

See, you're getting it! Now you have: \(\Large a^{-6}a^6b^2b^{12}\) Which you can mentally think of as grouped according to where you have the base of a, and where you have the base of b, e.g., think of it as: \(\Large (a^{-6}\cdot a^6)(b^2\cdot b^{12})\) And now you'll apply that last rule: \(\Large x^m\cdot x^n=x^{m+n}\) E.g. \(\Large 3^2\cdot 3^5=3^{7}\) \(\Large z^3\cdot z^8=z^{11}\) etc.... Now use that on: \(\Large (a^{-6}\cdot a^6)(b^2\cdot b^{12})\)

OpenStudy (debbieg):

(you'll use it trice, once on the "a" part and again on the "b" part)

OpenStudy (debbieg):

*twice lol

OpenStudy (anonymous):

\[a ^{0}b ^{14}\]

OpenStudy (debbieg):

Perfect! Now, do you know the rule for \(x^0\)? What is any base, raised to the power 0?

OpenStudy (anonymous):

1

OpenStudy (debbieg):

Exactly! So the fully simplified form of this expression is...?

OpenStudy (anonymous):

1\[1b ^{14}\]

OpenStudy (debbieg):

That's right! Now, you don't really need the "1"out in front... e.g..... "1x" is really just written as "x" in fully simplified form, follow me?

OpenStudy (anonymous):

yes thank you so much

OpenStudy (debbieg):

\(\begin{matrix}\Huge \color{gold}{★★★★★★★★★★}\\ \Huge \color{red}{\text{Great job!}} \\ \Huge \color{red}{\text{Gold stars for you!}}\\ \Huge \begin{matrix}\color{gold}{★★★}\\\color{gold}{★★}\\\color{gold}{★★★}\end{matrix} \color{purple}{\overbrace{\left(\begin{matrix}\color{blue}{\widehat\odot \widehat\odot} \\\small\bigtriangleup \\ \Huge \color{red}\smile\end{matrix}\right)}} \Huge \begin{matrix}\color{gold}{★★★}\\\color{gold}{★★}\\\color{gold}{★★★}\end{matrix}\end{matrix}\)

OpenStudy (debbieg):

no problem, happy to help. :)

OpenStudy (anonymous):

yay gold star